North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size

Wednesday 28th June 2017

A very wet and miserable day at Grosmont today, with Derek, Jon ,Ian Chris and Nigel. Not a great deal to report today, Nigel continuing painting the frames two coats of primer applied to each side. Derek has cleaned all the rubbish from the inside of the drag box and found a spoon and the missing metal ticket from the" Staff & Ticket" section Newbridge - Levisham ,which was lost about two years ago. Jon and I assembled the L/H valve together loosely keys still to make for these valves.

The dome stud hole was re- tapped and a new stud fitted - Ian Pearson

On the dome the stud hole was re- tapped and a new stud fitted see photo. There will be no working parties on Saturday 1/07/17, Wednesday, 5/07/17, Saturday 8/07/17, Wednesday 12/07/17. This is due to holidays and other commitments.